# FMIP: Joint Continuous-Integer Flow for Mixed-Integer Linear Programming

### v2

@@ 6 Conclusion
Appendix A Graph Representation Details
@@
  • Edges ($\mathcal{E}$): We construct a sparse bipartite graph based on the nonzero entries of the constraint matrix $\mathbf{A}$. Specifically, an edge is added between variable node $j$ and constraint node $i$ if and only if $\mathbf{A}[i,j]>0$. The edge feature is set as the corresponding coefficient, i.e., the edge from $j$ to $i$ carries weight $\mathbf{A}[i,j]$.
@@
Appendix B Graph Neural Networks
The input to our neural network is a tuple $(\boldsymbol{G}_{t},t),$ where $\boldsymbol{G}_{t}$ represents the solution graph as described in Section A. This graph encodes the relationships between variables and constraints, effectively capturing the structure of the MILP problem. The scalars $t$ and $t$ are temporal parameters associated with the integer and continuous variables, respectively, with both $t\in[0,1]$. These time parameters are introduced to enable a dynamic, progressive refinement of the solution space during the training process, aligning with the flow matching paradigm where time plays a critical role in guiding the evolution of the generated solution. We’ll details our neural networks structure in the following subsection.
B.1 TripartiteGCNConv Layer
@@ B.2 Model Structure
Message Passing Layers  For each layer $\ell=1,\dots,L$, we update the node representations via a message passing scheme over a tripartite graph. The relevant edge sets are defined as:
@@
TriConv is a tripartite message aggregation module that aggregates signals from both integer and continuous variable nodes into constraint nodes. BiConv has the same architecture as TriConv, but applies bipartite message passing from constraints back to one variable type.
@@
Appendix C Implementation Details
@@ C.2 Infrastructure
C.3 Downstream Solvers
Table 5 summarizes the hyperparameter settings used by each downstream solver across different problem instances. The meaning of each parameter (e.g., $K$, $\alpha$, $k_{0}$, $\Delta$, etc.) is consistent with the notation defined in Appendix D, where detailed descriptions of downstream solvers and their hyperparameters are provided. For a fair comparison, we use the same parameters for the downstream solvers when evaluating FMIP and all baselines.
C.4 FMIP
@@
Appendix D Downstream Solver Algorithm
@@ D.4 Apollo-MILP
Appendix E Pseudo code for FMIP Inference

### v3

@@ 6 Conclusion
Acknowledgment
Lin’s research is partially supported by the National Natural Science Foundation of China (NSFC) [Grant NSFC-624B2096, 72542012, 72595872]. Ge’s research is partially supported by the National Natural Science Foundation of China (NSFC) [Grant NSFC-72225009, 72394360, 72394365].
Appendix A LLM Usage Declaration
In this submission, we used a Large Language Model (LLM) solely for language refinement and text polishing. The LLM was employed to enhance the clarity, flow, and readability of the manuscript, but it did not contribute to the ideation, analysis, or generation of scientific content. All ideas, interpretations, and results presented in the paper are solely the work of the authors. No content generated by the LLM was used to fabricate facts or contribute to the research findings.
Appendix B Graph Representation Details
@@
  • Edges ($\mathcal{E}$): We construct a sparse bipartite graph based on the nonzero entries of the constraint matrix $\mathbf{A}$. Specifically, an edge is added between variable node $j$ and constraint node $i$ if and only if $\mathbf{A}[i,j]{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\not=}0$. The edge feature is set as the corresponding coefficient, i.e., the edge from $j$ to $i$ carries weight $\mathbf{A}[i,j]$.
@@
Appendix C Graph Neural Networks
The input to our neural network is a tuple $(\boldsymbol{G}_{t},t),$ where $\boldsymbol{G}_{t}$ represents the solution graph as described in Section B. This graph encodes the relationships between variables and constraints, effectively capturing the structure of the MILP problem. The scalars $t$ are temporal parameters associated with the integer and continuous variables, respectively, with both $t\in[0,1]$. These time parameters are introduced to enable a dynamic, progressive refinement of the solution space during the training process, aligning with the flow matching paradigm where time plays a critical role in guiding the evolution of the generated solution. We’ll details our neural networks structure in the following subsection.
C.1 GCNConv Layer
@@ C.2 Model Structure
Message Passing Layers  For each layer $\ell=1,\dots,L$, we update the node representations via a message passing scheme over a bipartite graph with three node types. The relevant edge sets are defined as:
@@
TriConv is a message aggregation module that aggregates signals from two types of source nodes into target nodes. BiConv has the same architecture as TriConv, but aggregates signals from one type of source nodes into target nodes.
@@
Appendix D Implementation Details
@@ D.2 Infrastructure
D.3 Downstream Solvers
Table 6 summarizes the hyperparameter settings used by each downstream solver across different problem instances. The meaning of each parameter (e.g., $K$, $\alpha$, $k_{0}$, $\Delta$, etc.) is consistent with the notation defined in Appendix E, where detailed descriptions of downstream solvers and their hyperparameters are provided. For a fair comparison, we use the same parameters for the downstream solvers when evaluating FMIP and all baselines.
D.4 FMIP
@@
Appendix E Downstream Solver Algorithm
@@ E.4 Apollo-MILP
Appendix F Pseudo code for FMIP Inference
@@
Appendix G Additional Experiments
G.1 Breakdown Analysis on MIPLIB 2017 Benchmark
To further investigate scalability and analyze failure cases, we partitioned the MIPLIB 2017 test set into five subsets based on the scale of the MILP instances. We report the relative primal gap, Rel.GAP, of FMIP on these subsets w.r.t. the Best Known Solution (BKS) provided by MIPLIB official website.
Table 8: Relative Primal Gap of FMIP on subsets of varying instance sizes.
Instance Size FMIP Rel. GAP Best Baseline Rel. GAP
50 – 3,000 4.70% 5.03%
3,000 – 8,500 4.81% 5.19%
8,500 – 20,000 4.86% 5.445%
20,000 – 70,000 9.04% 9.10%
70,000 – 300,000 23.26% 25.38%
The results in Table 8 reveal a distinct performance dichotomy: while FMIP maintains robust stability (Rel. GAP $5\%$) across small to medium-large instances (up to 20,000 variables), performance degrades on ultra-large-scale instances ($\text{70,000}$ variables). In these extreme regimes, the GAP widens to $\sim 23\%$ as the exponentially increasing complexity of the constraint landscape challenges the model’s generative precision.
Despite the degradation in extreme cases, FMIP still serves as the SOTA primal heuristic predictor to provide feasible starting points, making it significantly faster than a cold-start solver. This analysis provides a clear direction for future work, i.e., improving backbone scalability for ultra-large instances.
G.2 Scalability on Extremely Hard Instances
We further test FMIP combined with Predict-and-Search (FMIP-PS) against Gurobi on extremely large-scale Set Covering (SC) problems (up to 1,000,000 variables), which are significantly larger and harder than the training set (around 5,000 variables). The time limit of FMIP is set to 3600s, which is empirically enough for superior performance in such extreme cases. The time limit of Gurobi is set according to the problem scale.
Table 9: Objective Value over Time w.r.t. Different Test Problem Scale (Lower is Better).
(a) #Constraints = 50,000; #Variables = 500,000
Solver 60s 600s 1800s 3600s
Gurobi 785,972 785,972 53,008 52,877
FMIP-PS 1,335,583 52,984 52,905 52,859
(b) #Constraints = 100,000; #Variables = 500,000
Solver 60s 1800s 3600s 7200s
Gurobi 1,015,955 1,015,955 1,015,955 1,015,955
FMIP-PS 1,007,549 1,007,549 76,579 —Stop—
(c) #Constraints = 500,000; #Variables = 500,000
Solver 60s 1800s 3600s 7200s 25000s
Gurobi 1,654,768 818,547 469,130 469,130 469,130
FMIP-PS 1,007,549 184,821 184,821 —Stop— —Stop—
(d) #Constraints = 100,000; #Variables = 1,000,000
Solver 60s 1800s 3600s 7200s 25000s
Gurobi 1,022,627 1,022,627 1,022,627 1,022,627 44,039
FMIP-PS 192,417 192,417 43,113 —Stop— —Stop—
We can observe that, for these hard instances, Gurobi often stagnates or converges very slowly. FMIP could help Gurobi to find solutions that are orders of magnitude better (e.g., $\sim 43\text{k}$ vs $\sim 1\text{M}$) or reaches comparable solutions hours faster. These results further demonstrate the superiority of our proposed FMIP to further boost Gurobi, and validate the core practice of accelerating downstream solvers with AI powers.
Appendix H Relationships between FMIP and Exact Solvers like Gurobi
We would like to further clarify that FMIP is designed to enhance exact downstream solvers like Gurobi, instead of replacing them.
The exact solver, Gurobi, is the backend engine in our experiments. The “Downstream Solvers” in our paper (i.e., Neural Diving, Predict-and-Search, Apollo) are, in fact, neural decomposition frameworks that wrap the exact solver Gurobi as improved version. They do not replace Gurobi; rather, they use the heuristic solution predicted by FMIP to decompose the original hard MILP into smaller sub-MIPs, which are then solved by Gurobi. Therefore, our method is designed precisely to function with wrapped or exact solvers, enhancing their ability to find high-quality feasible solutions in restricted search spaces.